OpenStudy (anonymous):

if α and β are acute angles and cos2α=3cos2β-1/3-cos2β , show that tanα=√2 tanβ

OpenStudy (anonymous):

3-cos2β is the denominator for 3cos2β too....

OpenStudy (anonymous):

This is just manipulation of trigonometric identities.

OpenStudy (anonymous):

\[cos (2\alpha )= \dfrac{3cos (2\beta) -1}{3-cos(2\beta)}\]

OpenStudy (anonymous):

^^ Exactly

OpenStudy (anonymous):

wow, it's tough man!!

OpenStudy (anonymous):

I know.... :/

OpenStudy (anonymous):

@wio help!

OpenStudy (anonymous):

Surrender!! I don't know how to prove. I am sorry for wasting your time

myininaya (myininaya):

the thing you get to assume has double angles while think you want to show doesn't i think this is our starting point

OpenStudy (anonymous):

@myininaya we have formula link them together. It is \(\huge tan\alpha =\dfrac{1-cos2\alpha}{sin2\alpha}\) and I replace cos 2 alpha but I got nothing :(

OpenStudy (anonymous):

I tried many ways, even thought of to figure what it is. Nothing rings to me. :( 

myininaya (myininaya):

Right now what I'm doing is playing with it. If I come up with something I will give you a hint in that direction. Don't stop trying though. Just play and try to have fun. It is like a wonderful puzzle.

myininaya (myininaya):

i have something you are still here

myininaya (myininaya):

your triangle idea actually gave it to me

myininaya (myininaya):

except kinda

myininaya (myininaya):

so we are given \[\cos(2 \alpha)=\frac{3\cos(2 \beta)-1}{3-\cos(2 \beta)}\] I built a triangle with 2 alpha as a measurement for a right triangle. I use the top of that above ratio as the measurement for adjacent side of that triangle and the bottom of that ratio be the hyp measurement for that triangle

myininaya (myininaya):

then I found the opp side in terms of those measurements

myininaya (myininaya):

then you said \[\tan(\alpha)=\frac{1-\cos(2 \alpha)}{\sin(2 \alpha)}\] so I plug in what was given for cos(2 alpha) and what i found for sin(2 alpha)

myininaya (myininaya):

then get rid of the compound fraction

myininaya (myininaya):

and you will start to see something beautiful

myininaya (myininaya):

like it is more than beautiful is actually what you want to obtain you will have to use the fact that \[\tan(\beta)=\frac{1-\cos(2 \beta)}{\sin(2 \beta)}\] at the very end there

myininaya (myininaya):

ok well if you return let me know if you tried this and if you got stuck somewhere...

OpenStudy (mathmate):

@myininaya I am pretty close to getting it, but trying to find my arithmetical error. the denominator seems to reduce to cos^2(b), and the numerator seems to be edging close to cos^2(b)-2sin^2(b). I'll write it up and see if it makes sense.

myininaya (myininaya):

Oops. Sorry @bebong I messaged you when this was actually @bajji 's problem. My bad!

myininaya (myininaya):

Did you do my approach or another approach.

OpenStudy (mathmate):

@myininaya I just blindly expand everything and hope to put back together something. lol :)

myininaya (myininaya):

 given then I found opposite side: \[\text{ opp side measurement of } (2 \alpha)=\sqrt{(3-\cos(2 \beta))^2-(3\cos(2 \beta)-1)^2} \\ =\sqrt{(9-6\cos(2 \beta)+\cos^2(2 \beta))-(9\cos^2(2 \beta)-6\cos(2 \beta)+1) } \\ = \sqrt{8-8\cos^2(2 \beta)}=2 \sqrt{2} \sqrt{1-\cos^2(2 \beta)}\] Just a little more detail about the direction I took See how lovely that looks.